Twenty students have applied for merit scholarships. This year 5 merit scholarships were awarded. If a random sample of 5 applications (from the population of 20) is selected, what is the probability that at least 3 students were the recipients of scholarship? NOTE: WRITE YOUR ANSWER USING 4 DECIMAL DIGITS. FOR EXAMPLE 0.1234 OR 3.2450. DO NOT ROUND UP OR DOWN THE LAST DECIMAL DIGIT.
Answer #1

P9 No Answer:- Given that, N-20 M-5 n=5 Xw Hypergeometric CN = 20, M=5, n=5) PCX23) P(x=3) + P(x=4) +P(x=5) (36-3) +36-40)+1